a)How do I find the arithmetic mean of:
i) 9 and 21
ii) 6+sqrt(3) and 6-sqrt(3)
b) Insert 5 arithmetic means between 11 and 71.
Thank you very much for your assistance!

Answer by stanbon(75887)

a)How do I find the arithmetic mean of:
i) 9 and 21
---
Ans: (9+21)/2 = 10
------------------------
ii) 6+sqrt(3) and 6-sqrt(3)
Ans: Add and divide by 2 to get 6
----------------------------------
b) Insert 5 arithmetic means between 11 and 71.
a(1) = 5
a(6) = a(1) + 6d = 71
---
Solve for "d":
5 + 6d = 71
6d = 66
d = 11
---
Sequence: 5,16,27,38,49,60,71
======================================
Cheers,
Stan H.
=============